Q17:-1 . it takes 4.00 J of work to stretch a Hokes-law spring 100 cm fro m its unstressed length, determine the extra work required to stretch it an additional 5.0 cm. (A)6. 24) (B)7.56j (0)2.761 (D)3.84) (E)5.00
Answer #1

Work-done to stretch a spring is given by:

W = (1/2)k*x^2

given that

W = 4 J and x stretch = 10 cm = 0.1 m

So,

k = 2W/x^2

k = 2*4/0.1^2 = 800 N/m

Now to stretch it additionally 5 cm,

total stretch = 10 cm + 5 cm = 15 cm = 0.15 m

W = (1/2)k*x^2 = 800*0.15^2/2

W = 9 J

Extra work required = 9 J - 4 J = 5 J

Correct option is E.

Please Upvote.